# Consistency of Neural Causal Partial Identification

### v1

@@ 4.1 Experiments
In this section, we examine the performance of our algorithm in two settings. We compare our algorithm with the Autobounds algorithm [14] in a binary IV example in [14] and in a continuous IV model. Since Autobounds can only deal with discrete models, we discretize the continuous model for comparison purposes. The implementation details are provided in the Appendix D. The setting of the first experiment is taken from [14, Section D.1]. This is a binary IV problem and we can calculate the optimal bound using LP [4]. We find that our bound is close to the optimal bound. The second experiment is a general IV example where the treatment is binary but the rest of the variables are continuous. The details of the structural equations can be found in Appendix D. The program that Autobounds solves after discretization contains $\approx 2^{14}$ variables. Even with such a fine discretization, the bound obtained by Autobounds is not tighter than our NCM approach.
@@ Appendix A Illustration of Notions in Section 2
(b) The causal graph of this SCM.
@@
where $U_{1}$ and $U_{2}$ are correlated and $U_{3},U_{4}$ and $(U_{1},U_{2})$ are independent. The causal model is shown in Figure 2(a) and its causal graph is shown in Figure 2(b). In this example, $\boldsymbol{U}_{V_{1}}=\{U_{1}\},\boldsymbol{U}_{V_{2}}=\{U_{1},U_{2}\},\boldsymbol{U}_{V_{3}}=\{U_{2},U_{4}\},\boldsymbol{U}_{V_{4}}=\{U_{3},U_{4}\},\boldsymbol{U}_{V_{5}}=\{U_{3}\}$. Since $U_{1},U_{2}$ are correlated, $C_{1}=\{V_{1},V_{2},V_{3}\}$ is one $C^{2}$ component because all nodes in $C_{1}$ are connected by bi-directed edges. Note that $\{V_{1},V_{2},V_{3},V_{4}\}$ is not a $C^{2}$ component because $V_{4}$ and $V_{2}$ is not connected by any bi-directive edge. The rest of $C^{2}$ components are $C_{2}=\{V_{4},V_{5}\},C_{3}=\{V_{3},V_{4}\}$.
@@ Appendix B Proof of approximation Theorems
B.1 Proof of Theorem 1
@@
Suppose that 1 and 2 hold, given any SCM $\mathcal{M}$ with causal graph $\mathcal{G}$ and latent variables $\boldsymbol{U}$, we can construct a canonical SCM $\hat{\mathcal{M}}$ of the form (3) by merging the latent variables in $\boldsymbol{U}$ such that $\mathcal{M}$ and $\hat{\mathcal{M}}$ produce the same intervention results. Besides, functions in $\hat{\mathcal{M}}$ have the same smoothness as $\mathcal{M}$.
Proof of Proposition 4.
@@
Let $\boldsymbol{G}=\left\{G_{V_{i}}:V_{i}\ \text{is categorical}\right\}$ and the latent variables in the original model $\mathcal{M}$ be $\boldsymbol{U}=\{U_{1},\cdots,U_{n_{U}}\}\cup\boldsymbol{G}$. By 1 and 2, $\boldsymbol{U}$ are independent and the structure equations of $\mathcal{M}$ have the form
$V_{i}=\begin{cases}f_{i}\left(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}^{V_{i}}+\log\left(f_{i}\left(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\end{cases}$
@@
where $g_{k}^{V_{i}}$ are i.i.d. standard Gumbel variables and $\boldsymbol{U}_{V_{i}}\subset\{U_{1},\cdots,U_{n_{U}}\}$ contains the latent variables that affect $V_{i}$. We regroup and merge the variables ${U}_{i}$ to make the model have a canonical form while not changing the functions in the structure equations. Let $D_{1},\cdots,D_{n_{C}}\subset\boldsymbol{V}$ be the $C^{2}$-component of $\mathcal{M}$. For each ${U}_{i}$, we define the vertices that are affected by $U_{i}$ as
$I({U}_{i})=\{V_{j}:{U}_{i}\in{\boldsymbol{U}}_{V_{j}}\}.$
@@
We partition $\{U_{1},\cdots,U_{n_{U}}\}$ into $n_{C}$ sets $\hat{\boldsymbol{U}}_{1},\cdots,\hat{\boldsymbol{U}}_{n_{C}}$ in the following way. For each ${U}_{k}$, ${U}_{k}$ is in the set $\hat{\boldsymbol{U}}_{i}$ if $I({U}_{k})\subset D_{i}$. If there are two components $D_{i},D_{j}$ satisfy the condition, we put ${U}_{k}$ into either of the sets $\hat{\boldsymbol{U}}_{i},\hat{\boldsymbol{U}}_{j}$. Let $\hat{U}_{i}$ have the same distribution as the joint distribution of the random variables in set $\hat{\boldsymbol{U}}_{i}$. Let ^$\hat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},{\mathcal{F}},\hat{\boldsymbol{U}})$ the SCM with structure equations
@@
$\hat{V}_{i}=\begin{cases}f_{i}\left(\text{Pa}(\hat{V}_{i}),\hat{U}_{k_{1}},\cdots,\hat{U}_{k_{n_{i}}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}^{V_{i}}+\log\left(f_{i}\left(\text{Pa}(\hat{V}_{i}),\hat{U}_{k_{1}},\cdots,\hat{U}_{k_{n_{i}}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\end{cases},\hat{\boldsymbol{U}}_{k_{i}}\cap\boldsymbol{U}_{V_{i}}\neq\emptyset,\ i\in[n_{V}].$
@@
Here, we slightly abuse the notation ${f}_{i}$. ${f}_{i}$ ignores inputs from $\hat{\boldsymbol{U}}_{k_{1}},\cdots,\hat{\boldsymbol{U}}_{k_{n_{i}}}$ that are not in ${\boldsymbol{U}_{V_{i}}}$. Note that $\hat{\boldsymbol{U}}_{1},\cdots,\hat{\boldsymbol{U}}_{n_{C}},\boldsymbol{G}$ has the same distribution as ${\boldsymbol{U}}$ because we only merge some latent variables. In addition, the functions in the new model $\hat{\mathcal{M}}$ has the same smoothness as the original model.
@@
We first verify that the causal graph of $\hat{\mathcal{M}}$ is $\mathcal{G}$. If there is a bi-directed arrow between nodes $V_{i},V_{j}$ in $\mathcal{G}$, by the independence assumption, there must be one latent variable ${U}_{k}\in{\boldsymbol{U}}_{V_{i}}\cap{\boldsymbol{U}}_{V_{j}}$. There exist one $D_{l}$ such that $I({U}_{k})\subset D_{l}$ and $\hat{\boldsymbol{U}}_{l}\cap{\boldsymbol{U}}_{V_{i}}\neq\emptyset,\hat{\boldsymbol{U}}_{l}\cap{\boldsymbol{U}}_{V_{j}}\neq\emptyset$. Therefore, $\hat{U}_{l}$ will affect both $V_{i}$ and $V_{j}$ and there is a bi-directed arrow between node $V_{i},V_{j}$ in $\mathcal{G}_{\hat{\mathcal{M}}}$. Suppose that there is a bi-directed arrow between node $V_{i},V_{j}$ in $\mathcal{G}_{\hat{\mathcal{M}}}$, it means there exist a $\hat{U_{k}}$ such that $\hat{\boldsymbol{U}}_{k}\cap{\boldsymbol{U}}_{V_{i}}\neq\emptyset,\hat{\boldsymbol{U}}_{k}\cap{\boldsymbol{U}}_{V_{j}}\neq\emptyset$. Let ${U}_{l_{1}}\in\hat{\boldsymbol{U}}_{k}\cap{\boldsymbol{U}}_{V_{i}},{U}_{l_{2}}\in\hat{\boldsymbol{U}}_{k}\cap{\boldsymbol{U}}_{V_{j}}$, then
@@
Since $D_{k}$ is a $C^{2}$-component, there exist a bi-directed arrow between node $V_{i},V_{j}$ in $\mathcal{G}$. Therefore, $\mathcal{G}=\mathcal{G}_{\hat{\mathcal{M}}}$.
Finally, we verify that $\mathcal{M}$ and $\hat{\mathcal{M}}$ produce the same intervention results. The intervention distribution can be viewed as the push-forward of the latent distribution, i.e., $\boldsymbol{V}=F(\boldsymbol{U})$, and the intervention operation only changes the function $F$. In our construction, we only merge the latent variables and leave the functions in structure equations being the same (except that they may ignore some coordinates in input). For any intervention $T=t$, suppose in $\mathcal{M}$ we have $\boldsymbol{V}(t)=F_{t}(\boldsymbol{U})$. Then, in $\hat{\mathcal{M}}$, we get $\hat{\boldsymbol{V}}(t)=F_{t}(\hat{\boldsymbol{U}})$. Since $\boldsymbol{U}$ and $\hat{\boldsymbol{U}}$ has the same distribution, the distribution of ${\boldsymbol{V}}(t),\hat{\boldsymbol{V}}(t)$ are the same. ∎
Proof of Theorem 1.
@@
$V_{i}=\begin{cases}f_{i}\left(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}^{V_{i}}+\log\left(f_{i}\left(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\|f_{i}\|_{1}=1,\end{cases}$
where $g_{k}^{V_{i}}$ are i.i.d. standard Gumbel variables and ^$\hat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},\hat{\mathcal{F}},\hat{\boldsymbol{U}})$ has structure equations
$\hat{V}_{i}=\begin{cases}\hat{f}_{i}\left(\text{Pa}(\hat{V}_{i}),\hat{\boldsymbol{U}}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{\hat{g}_{k}^{V_{i}}+\log\left(\hat{f}_{i}\left(\text{Pa}(\hat{V}_{i}),\hat{\boldsymbol{U}}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\|\hat{f}_{1}\|_{1}=1,\end{cases}$
where $\hat{g}_{k}^{V_{i}}$ are i.i.d. standard Gumbel variables. Let the treatment variables set be $\boldsymbol{T}=\{T_{1},\cdots,T_{n_{T}}\}$. We may give all vertices and $\boldsymbol{U}_{0}=\{U_{1},\cdots,U_{n_{U}}\}$, an topology ordering (rearrange the subscript if necessary)
$U_{1},\cdots,U_{n_{U}},T_{1},\cdots,T_{n_{T}},V_{1}(\boldsymbol{t}),\cdots,V_{n_{V}-n_{T}}(\boldsymbol{t})$
such that for each directed edge $(V_{i}(\boldsymbol{t}),V_{j}(\boldsymbol{t}))$, vertex $V_{i}(\boldsymbol{t})$ lies before vertex $V_{j}(\boldsymbol{t})$ in the ordering, ensuring that all edges start from vertices that appear early in the order and end at vertices that appear later in the order. We put $\boldsymbol{U}_{0}$ and $\boldsymbol{T}$ at the beginning because they are the root nodes of the intervened model. We denote $\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k}^{\boldsymbol{T}}$ (resp. $\hat{\mu}_{\boldsymbol{U}_{0},\boldsymbol{T},1:k}^{\boldsymbol{T}}$) to be the distribution of $\boldsymbol{U}_{0},\boldsymbol{T},V_{1}(\boldsymbol{t}),\cdots,V_{k}(\boldsymbol{t})$ (resp. $\hat{\boldsymbol{U}}_{0},\boldsymbol{T},\hat{V}_{1}(\boldsymbol{t}),\cdots,\hat{V}_{k}(\boldsymbol{t})$), $\mu_{V_{k}|\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}$ the distribution of $V_{k}$ given $\boldsymbol{U}_{0},\boldsymbol{T},V_{1}(\boldsymbol{t}),\cdots,V_{k-1}(\boldsymbol{t})$.
Let $S=\sum_{i=1}^{n_{V}}\|f_{i}-\hat{f}_{i}\|_{\infty}$. Next, we prove that
$W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k}^{\boldsymbol{T}}\right)\leqslant(L+1)W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right)+2K^{2}S.$ (8)
@@
$W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k}^{\boldsymbol{T}}\right)$ $=\sup_{g\in Lip(1)}\int g(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k})d\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k}^{\boldsymbol{T}}-\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k}^{\boldsymbol{T}}\right)$

### v2

@@ 4.1 Experiments
In this section, we examine the performance of our algorithm in two settings. We compare our algorithm with the Autobounds algorithm [14] in a binary IV example in [14] and in a continuous IV model  [^4] The code can be found in https://github.com/Jiyuan-Tan/NeuralPartialID. Since Autobounds can only deal with discrete models, we discretize the continuous model for comparison purposes. The implementation details are provided in the Appendix D. The setting of the first experiment is taken from [14, Section D.1]. This is a binary IV problem and we can calculate the optimal bound using LP [4]. We find that our bound is close to the optimal bound. The second experiment is a general IV example where the treatment is binary but the rest of the variables are continuous. The program that Autobounds solves after discretization contains $\approx 2^{14}$ variables. Even with such a fine discretization, the bound obtained by Autobounds is not tighter than our NCM approach. The details of the structural equations and analysis can be found in Appendix D. We also provide an extra experiment on the counterexample of Proposition C.1 in the appendix to show the effect of Lipschitz regularization.
@@
Conclusion
In this paper, we provide theoretical justification for using NCMs for partial identification. We show that NCMs can be used to represent SCMs with complex unknown latent distributions under mild assumptions and prove the asymptotic consistency of the max/min estimator for partial identification of causal effects in general settings with both discrete and continuous variables. Our results also provide guidelines on the practical implementation of this method and on what hyperparameters are important, as well as recommendations on values that these hyperparameters should take for the consistency of the method. These practical guidelines were validated with a small set of targeted experiments, which also showcase superior performance of the neural-causal approach as compared to a prior main contender approach from econometrics and statistics, that involves discretization and polynomial programming.
An obvious next step in the theoretical foundation of neural-causal models is providing finite sample guarantees for this method, which requires substantial further theoretical developments in the understanding of the geometry of the optimization program that defines the bounds on the causal effect of interest. We take a first step in that direction for the special case, when there are no unobserved confounders and view the general case as an exciting avenue for future work.
Acknowledgement
Vasilis Syrgkanis is supported by NSF Award IIS-2337916 and a 2022 Amazon Research Award.
@@ Appendix A Illustration of Notions in Section 2
Figure 4: The causal graph of this SCM.
@@
where $U_{1}$ and $U_{2}$ are correlated and $U_{3},U_{4}$ and $(U_{1},U_{2})$ are independent. The causal model is shown in Figure 4 and its causal graph is shown in Figure 4. In this example, $\bm{U}_{V_{1}}=\{U_{1}\},\bm{U}_{V_{2}}=\{U_{1},U_{2}\},\bm{U}_{V_{3}}=\{U_{2},U_{4}\},\bm{U}_{V_{4}}=\{U_{3},U_{4}\},\bm{U}_{V_{5}}=\{U_{3}\}$. Since $U_{1},U_{2}$ are correlated, $C_{1}=\{V_{1},V_{2},V_{3}\}$ is one $C^{2}$ component because all nodes in $C_{1}$ are connected by bi-directed edges. Note that $\{V_{1},V_{2},V_{3},V_{4}\}$ is not a $C^{2}$ component because $V_{4}$ and $V_{2}$ is not connected by any bi-directive edge. The rest of $C^{2}$ components are $C_{2}=\{V_{4},V_{5}\},C_{3}=\{V_{3},V_{4}\}$.
@@ Appendix B Proof of approximation Theorems
B.1 Illustration of how to construct canonical representations and neural architectures
In this section, we illustrate how to construct canonical representations and neural architectures given a causal graph via a simple example. We consider the example in the previous section. The causal graph is shown in Figure 4.
[…]
Figure 6: The canonical representation of Figure 4.
[…]
Figure 7: The neural network architecture.
Following the construction in Proposition B.1, we use one latent variable for each $C^{2}$ component. As we explain in Appendix A, this causal model has three $C^{2}$ components, $\{V_{1},V_{2},V_{3}\},\{V_{3},V_{4}\},\{V_{4},V_{5}\}$. In the canonical representation, exactly the latent variables enter their corresponding $C^{2}$ component. The structure equation of this SCM is as follows.
$V_{1}=f_{1}(V_{2},E_{1}),$
$V_{2}=f_{2}(E_{1}),$
$V_{3}=f_{3}(V_{1},V_{2},E_{1},E_{2}),$ (B.1)
$V_{4}=f_{4}(V_{3}.E_{2}),$
$V_{5}=f_{5}(V_{4},E_{3}),$
If we set $E_{1}=(U_{1},U_{2}),E_{2}=U_{4},E_{3}=U_{3}$, (B.1) is equivalent to (A.1). Therefore, we can see from this example that the canonical representation does not lose any information about the SCM.
Now, we show how to construct the NCM architecture from a canonical representation. As we mentioned in Section 3, we approximate the latent distribution by pushing forward uniform and Gumbel variables. The structure equation of the NCM is
$V_{1}=f^{\theta_{1}}_{1}(V_{2},g_{1}^{\theta_{1}}(Z_{1})),$
$V_{2}=f^{\theta_{2}}_{2}(g_{1}^{\theta_{1}}(Z_{1})),$
$V_{3}=f^{\theta_{2}}_{3}(V_{1},V_{2},g_{1}^{\theta_{1}}(Z_{1}),g_{2}^{\theta_{1}}(Z_{2})),$ (B.2)
$V_{4}=f_{4}^{\theta_{4}}(V_{3}.g_{2}^{\theta_{1}}(Z_{2})),$
$V_{5}=f_{5}^{\theta_{5}}(V_{4},g_{2}^{\theta_{3}}(Z_{3})),$
where $f_{i}^{\theta_{i}},g_{j}^{\theta_{j}}$ are neural networks, $Z_{i}$ are join distribution of independent uniform and Gumbel variables, and $g_{j}^{\theta_{j}}$ has the special architecture described in Section 3.1 and Section 3.2. Figure 7 shows the architecture of the NCM. Each in-edge represents an input of a neural net.
B.2 Proof of Theorem 3.3
@@ B.1 Illustration of how to construct canonical representations and neural architectures
Suppose that 2.3 and 2.4 hold, given any SCM $\mathcal{M}$ with causal graph $\mathcal{G}$ and latent variables $\bm{U}$, we can construct a canonical SCM $\widehat{\mathcal{M}}$ of the form (3.1) by merging the latent variables in $\bm{U}$ such that $\mathcal{M}$ and $\widehat{\mathcal{M}}$ produce the same intervention results. Besides, functions in $\widehat{\mathcal{M}}$ have the same smoothness as $\mathcal{M}$.
Proof of Proposition B.1.
@@
Let $\bm{G}=\left\{G_{V_{i}}:V_{i}\ \text{is categorical}\right\}$ and the latent variables in the original model $\mathcal{M}$ be $\bm{U}=\{U_{1},\cdots,U_{n_{U}}\}\cup\bm{G}$. By 2.3 and 2.4, $\bm{U}$ are independent and the structure equations of $\mathcal{M}$ have the form
$V_{i}=\begin{cases}f_{i}\left(\text{Pa}(V_{i}),\bm{U}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}^{V_{i}}+\log\left(f_{i}\left(\text{Pa}(V_{i}),\bm{U}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\end{cases}$
@@
where $g_{k}^{V_{i}}$ are i.i.d. standard Gumbel variables and $\bm{U}_{V_{i}}\subset\{U_{1},\cdots,U_{n_{U}}\}$ contains the latent variables that affect $V_{i}$. We regroup and merge the variables ${U}_{i}$ to make the model have a canonical form while not changing the functions in the structure equations. Let $D_{1},\cdots,D_{n_{C}}\subset\bm{V}$ be the $C^{2}$-component of $\mathcal{M}$. For each ${U}_{i}$, we define the vertices that are affected by $U_{i}$ as
$I({U}_{i})=\{V_{j}:{U}_{i}\in{\bm{U}}_{V_{j}}\}.$
@@
We partition $\{U_{1},\cdots,U_{n_{U}}\}$ into $n_{C}$ sets $\widehat{\bm{U}}_{1},\cdots,\widehat{\bm{U}}_{n_{C}}$ in the following way. For each ${U}_{k}$, ${U}_{k}$ is in the set $\widehat{\bm{U}}_{i}$ if $I({U}_{k})\subset D_{i}$. If there are two components $D_{i},D_{j}$ satisfy the condition, we put ${U}_{k}$ into either of the sets $\widehat{\bm{U}}_{i},\widehat{\bm{U}}_{j}$. Let $\widehat{U}_{i}$ have the same distribution as the joint distribution of the random variables in set $\widehat{\bm{U}}_{i}$. Let ^$\widehat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},{\mathcal{F}},\widehat{\bm{U}})$ the SCM with structure equations
@@
$\widehat{V}_{i}=\begin{cases}f_{i}\left(\text{Pa}(\widehat{V}_{i}),\widehat{U}_{k_{1}},\cdots,\widehat{U}_{k_{n_{i}}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}^{V_{i}}+\log\left(f_{i}\left(\text{Pa}(\widehat{V}_{i}),\widehat{U}_{k_{1}},\cdots,\widehat{U}_{k_{n_{i}}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\end{cases},\widehat{\bm{U}}_{k_{i}}\cap\bm{U}_{V_{i}}\neq\emptyset,\ i\in[n_{V}].$
@@
Here, we slightly abuse the notation ${f}_{i}$. ${f}_{i}$ ignores inputs from $\widehat{\bm{U}}_{k_{1}},\cdots,\widehat{\bm{U}}_{k_{n_{i}}}$ that are not in ${\bm{U}_{V_{i}}}$. Note that $\widehat{\bm{U}}_{1},\cdots,\widehat{\bm{U}}_{n_{C}},\bm{G}$ has the same distribution as ${\bm{U}}$ because we only merge some latent variables. In addition, the functions in the new model $\widehat{\mathcal{M}}$ has the same smoothness as the original model.
@@
We first verify that the causal graph of $\widehat{\mathcal{M}}$ is $\mathcal{G}$. If there is a bi-directed arrow between nodes $V_{i},V_{j}$ in $\mathcal{G}$, by the independence assumption, there must be one latent variable ${U}_{k}\in{\bm{U}}_{V_{i}}\cap{\bm{U}}_{V_{j}}$. There exist one $D_{l}$ such that $I({U}_{k})\subset D_{l}$ and $\widehat{\bm{U}}_{l}\cap{\bm{U}}_{V_{i}}\neq\emptyset,\widehat{\bm{U}}_{l}\cap{\bm{U}}_{V_{j}}\neq\emptyset$. Therefore, $\widehat{U}_{l}$ will affect both $V_{i}$ and $V_{j}$ and there is a bi-directed arrow between node $V_{i},V_{j}$ in $\mathcal{G}_{\widehat{\mathcal{M}}}$. Suppose that there is a bi-directed arrow between node $V_{i},V_{j}$ in $\mathcal{G}_{\widehat{\mathcal{M}}}$, it means there exist a $\widehat{U_{k}}$ such that $\widehat{\bm{U}}_{k}\cap{\bm{U}}_{V_{i}}\neq\emptyset,\widehat{\bm{U}}_{k}\cap{\bm{U}}_{V_{j}}\neq\emptyset$. Let ${U}_{l_{1}}\in\widehat{\bm{U}}_{k}\cap{\bm{U}}_{V_{i}},{U}_{l_{2}}\in\widehat{\bm{U}}_{k}\cap{\bm{U}}_{V_{j}}$, then
@@
Since $D_{k}$ is a $C^{2}$-component, there exist a bi-directed arrow between node $V_{i},V_{j}$ in $\mathcal{G}$. Therefore, $\mathcal{G}=\mathcal{G}_{\widehat{\mathcal{M}}}$.
Finally, we verify that $\mathcal{M}$ and $\widehat{\mathcal{M}}$ produce the same intervention results. The intervention distribution can be viewed as the push-forward of the latent distribution, i.e., $\bm{V}=F(\bm{U})$, and the intervention operation only changes the function $F$. In our construction, we only merge the latent variables and leave the functions in structure equations being the same (except that they may ignore some coordinates in input). For any intervention $T=t$, suppose in $\mathcal{M}$ we have $\bm{V}(t)=F_{t}(\bm{U})$. Then, in $\widehat{\mathcal{M}}$, we get $\widehat{\bm{V}}(t)=F_{t}(\widehat{\bm{U}})$. Since $\bm{U}$ and $\widehat{\bm{U}}$ has the same distribution, the distribution of ${\bm{V}}(t),\widehat{\bm{V}}(t)$ are the same. ∎
Proof of Theorem 3.3.
@@
$V_{i}=\begin{cases}f_{i}\left(\text{Pa}(V_{i}),\bm{U}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}^{V_{i}}+\log\left(f_{i}\left(\text{Pa}(V_{i}),\bm{U}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\|f_{i}\|_{1}=1,\end{cases}$
where $g_{k}^{V_{i}}$ are i.i.d. standard Gumbel variables and ^$\widehat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},\widehat{\mathcal{F}},\widehat{\bm{U}})$ has structure equations
$\widehat{V}_{i}=\begin{cases}\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),\widehat{\bm{U}}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{\widehat{g}_{k}^{V_{i}}+\log\left(\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),\widehat{\bm{U}}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\|\widehat{f}_{1}\|_{1}=1,\end{cases}$
where $\widehat{g}_{k}^{V_{i}}$ are i.i.d. standard Gumbel variables. Let the treatment variables set be $\bm{T}=\{T_{1},\cdots,T_{n_{T}}\}$. We may give all vertices and $\bm{U}_{0}=\{U_{1},\cdots,U_{n_{U}}\}$, an topology ordering (rearrange the subscript if necessary)
$U_{1},\cdots,U_{n_{U}},T_{1},\cdots,T_{n_{T}},V_{1}(\bm{t}),\cdots,V_{n_{V}-n_{T}}(\bm{t})$
such that for each directed edge $(V_{i}(\bm{t}),V_{j}(\bm{t}))$, vertex $V_{i}(\bm{t})$ lies before vertex $V_{j}(\bm{t})$ in the ordering, ensuring that all edges start from vertices that appear early in the order and end at vertices that appear later in the order. We put $\bm{U}_{0}$ and $\bm{T}$ at the beginning because they are the root nodes of the intervened model. We denote $\mu_{\bm{U}_{0},\bm{T},1:k}^{\bm{T}}$ (resp. $\widehat{\mu}_{\bm{U}_{0},\bm{T},1:k}^{\bm{T}}$) to be the distribution of $\bm{U}_{0},\bm{T},V_{1}(\bm{t}),\cdots,V_{k}(\bm{t})$ (resp. $\widehat{\bm{U}}_{0},\bm{T},\widehat{V}_{1}(\bm{t}),\cdots,\widehat{V}_{k}(\bm{t})$), $\mu_{V_{k}|\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}}$ the distribution of $V_{k}$ given $\bm{U}_{0},\bm{T},V_{1}(\bm{t}),\cdots,V_{k-1}(\bm{t})$.
Let $S=\sum_{i=1}^{n_{V}}\|f_{i}-\widehat{f}_{i}\|_{\infty}$. Next, we prove that
$W\left(\mu_{\bm{U}_{0},\bm{T},1:k}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:k}^{\bm{T}}\right)\leqslant(L+1)W\left(\mu_{\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:k-1}^{\bm{T}}\right)+2K^{2}S.$ (B.3)
@@
$W\left(\mu_{\bm{U}_{0},\bm{T},1:k}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:k}^{\bm{T}}\right)$ $=\sup_{g\in Lip(1)}\int g(\bm{u},\bm{t},v_{1},\cdots,v_{k})d\left(\mu_{\bm{U}_{0},\bm{T},1:k}^{\bm{T}}-\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:k}^{\bm{T}}\right)$